Stock in Country Road Industries has a beta of 1.13. The market risk premium is 7.5 percent, and T-bills are currently yielding 4 percent. The company's most recent dividend was $1.6 per share, and dividends are expected to grow at a 5.5 percent annual rate indefinitely. If the stock sells for $31 per share, what is your best estimate of the company's cost of equity? (Do not round your intermediate calculations.)

Multiple Choice

  • 12.48%

  • 11.71%

  • 10.95%

  • 7.96%

  • 9.65%
Answer #1

cost of equity as per growth model=(D1/Current price)+Growth rate

=[(1.6*1.055)/31]+0.055=10.9451613%

cost of equity as per CAPM=risk free rate+Beta*market risk premium

=4+(1.13*7.5)=12.475%

Best estimate of  cost of equity=(10.9451613+12.475)/2

=11.71%(Approx).